OpenStudy (anonymous):

One isotope of nitrogen has 7 protons and 8 neutrons. Which is the correct reference for this isotope? A. nitrogen-8 B. nitrogen-7 C. nitrogen-15 D. nitrogen-1

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@izzy529

OpenStudy (anonymous):

Do you have any guesses?

OpenStudy (anonymous):

no

OpenStudy (gamer56):

@thatawesomebguy What do you think it cant be ?

OpenStudy (michele_laino):

I think nitrogen-15, since, we have to refer to the mass number A, which is the sum between the atomic number and the number of the neutrons inside the nucleus, so we have: A=Z+N= 7+ 8= 15